OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 4y = 2x + 10 8x − 3y = -14 a) (-1, 2) b) (7, 6) c) (-3, 2) d) (5, 5)

OpenStudy (anonymous):

since (-1, 2) satisfies both the given equations. therefore (-1, 2) is the corect ordered pair.

OpenStudy (anonymous):

can you help me with this one too Which of the following systems of equations has no solution? 9x + 5y = 1; 15y = 18x − 4 -7x − 7 = 3y; -14y − 8 = -6x 4x − 3y = 9; 6y = 8x − 18 7y = 5x − 10; 10x − 14y = 8

OpenStudy (anonymous):

4x − 3y = 9; 6y = 8x − 18

OpenStudy (anonymous):

can you help me on a few more please

OpenStudy (anonymous):

just going to have dinner please

OpenStudy (anonymous):

Select the ordered pair from the choices below that satisfies following system of equations: 2x − 10y = -14 -4y = -x − 5 (-1/2,1) (-20, 3) (3, 2) (8, -4)

OpenStudy (anonymous):

@AlwaystheBookworm

OpenStudy (anonymous):

What would like help with?

OpenStudy (anonymous):

help me with the question i just put the 2x − 10y = -14

OpenStudy (anonymous):

2x − 10y = -14 -4y = -x − 5 Alrighty, first change the equations to x+y=b 2x-10y=-14 x-4y=-5 Multiply the second equation by -2. 2x-10y=-14 -2x+8y=10 Cancel the x, add the ys and constants. -2y=-4 Divide by -2. y=2 Take the y and plug it into an original equation to find x.

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

x=3